# Received August 1, 2010; revised October 1, 2010; accepted for publication November 1, 2010

### v1

@@ 2.4 Expected value of sample information (EVSI)
Assume for the moment that we know the true case-mix ($\eta$) and the correct model ($\theta$). As such, we can generate data of the future study $D^{*}$ (Heath and others, 2021). Given the assumption on the exchangability of $D$ and $D^{*}$, $D^{*}$ would be combined with the $D$, which we denote by $D^{+}=(D,D^{*})$, and the investigator would revisit their decision for declaring the superior decision based on $W(D^{+})$. The ENB associated with a given $D^{+}$ is $ENB\left(W(D^{+});D^{+}\right)$.
However, as the future data have not been collected, we will need to take expectation with respect to the predictive distribution of $D^{*}$:
$EVSI=$ $E_{D^{+}}\left[ENB\left(W(D^{+});D^{+}\right)\right]-ENB(W(D);D)$ (9)
EVSI is a scalar quantity in the same NB unit as the clinical utility of the model. The higher the EVSI, the higher the expected benefit from a sample. The EVSI asymptotes to EVPI.
@@ 2.5 Bootstrap-based EVSI computation
The second term on the EVSI formula in (9) is the ENB and can be estimated following the same generic bootstrap algorithm given in Section 2.3. Therefore, to compute EVSI, we need to estimate the first term. As the equation for EVSI implies, the nested computation of the two expectations generally gives rises to nested Monte Carlo algorithms (Ades and others, 2004b; Kunst and others, 2020). Our non-parametric specification of $\eta$ gives rise to a nested algorithm based on the Bayesian bootstrap. We note that, because the Bayesian bootstrap allows one to interpret $D^{t}$ as a random draw from the population that has generated $d$, we can use $D^{t}$ to also generate the sample for the planned future study. We can generate a random draw from $D^{*}$ through another level of sampling. Specifically, let $m$ be the size of the future sample. We then take a (second level of) bootstrap sample from $D^{(t)}$, called $D^{(*t)}=(X^{(*t)},Y^{(*t)})$, as our future sample of size $n^{*}$ given $\theta^{(t)}$ (and consequently $\pi^{(t)}$). Note that, since we drew $D^{(*t)}$ under the assumption of $\pi^{(t)}$s being the true risks for the population individuals, we generate $Y^{(*t)}_{i}$s from Bernoulli distributions with their corresponding true risks being the success probabilities. Next, we update the original prediction risk model based on the pooled dataset of $D^{(+t)}=(D,D^{(*t)})$ as well as the predicted risk vector of $\pi\left(X^{(+t)},\hat{\theta}^{(t)}\right)$ obtained from the updated model. The bootstrap-based EVSI computation algorithm presented in Algorithm 1.
@@
  for $t=1,\dots,T$ do
     A random draw from the generating population $D^{(t)}$: bootstrap sample from original dataset
     True risks $\mathbf{\pi}^{(t)}$: estimated risks from the refitted model on $D^{(t)}$
     True NBs: estimate according to (5) on the development sample
     True NB of perfect information: estimate according to (5) where substitute $\pi\left(x_{j},\hat{(}\theta)\right)$ by its true value  [^1] See Sadatsafavi and others (2023b) for more details
     Future sample $D^{(*t)}$: bootstrap sample of size $m$ from $D^{(t)}$
     true outcomes: replace outcomes in $D^{(*t)}$ by a generated set of $Y^{(*t)}$ based on $\pi^{(t)}$
     Merged sample $D^{(+t)}$: merge the future sample $D^{(*t)}$ with the development sample $D$
     Updated risks $\pi\left(X^{(+t)},\hat{\theta}^{(t)}\right)$: estimated risks from the refitted model on $D^{(+t)}$
     Estimate NBs: estimate according to (5) on the merged sample
  end for
  Estimated ENBs: average estimated NBs based on $D$ according to (8)
  Estimated ENBs: average estimated NBs based on $D^{+}$s according to (8)
  Winner strategy: identify the strategy with maximum estimated ENB obtained from the previous step
  EVPI: Compute as the difference between ENB of perfect information and ENB of winner strategy
  EVSI: Compute according to (9)
@@ 3 Case study: GUSTO trial
As a case study, we apply our findings to GUSTO-I, a clinical trial of multiple thrombolytic strategies for acute myocardial infarction (AMI, or ’heart attack’) (investigators, 1993). The GUSTO trial was a four-arm randomized clinical trial that compared the efficacy of different intravenous thrombolytic regimens for AMI. This dataset has been widely used for methodological research in predictive analytics (Ennis and others, 1998). We use these data to build a prediction model for predicting 30-day mortality for AMI. We used sex, age, diabetes (yes or no), milocc (which part of the heart muscle is affected by the blocked artery and coded as 3=anterior, 4=inferior, 5=other, and 6=no MI), previous MI, hypertension, smoking, killip class, and treatment as covariates in the model. We consider risk threshold of 2% as the default threshold, but will evaluate the model at the entire possible range of thresholds.
Given the relatively large sample size (n=40830), the uncertainty in predictions when the model is fitted using the full sample is negligible. Indeed, a previous study reported an EVPI of 0 with the full dataset (thus, EVSI will be zero for all sample sizes [Sadatsafavi and others, 2023a]). In a similar fashion as Sadatsafavi and others (2023a), we took, without replacement, a random sub-sample of 1000 patients from GUSTO data and treat that as our development sample (current information).
@@
We start by drawing the decision curve at the entire possible range of thresholds, as shown in Figure 1 (panel A). The figure shows the NB of the model as well as the alternative default strategies (treat none and treat all). The model had higher expected NB than the alternative strategies at the all threshold values of interest. Additionally, panel (B) of Figure 1 shows the difference between the estimated NB of the model and that of the best alternative default strategy, i.e., $ENB(1,d)-max\left(0,ENB(2,d)\right)$, along with its bootstrap $95\%$ CI. Again, the model had a higher estimated NB that the default alternatives at the entire range of threshold values of interest.
Next, we draw EVSI. We chose four different future sample sizes: 100, 1000, 5000, and 10000. Following algorithm 1, for each sample size, we drew $10^{5}$ bootstrap samples ($T=10^{5}$) from the development sample, and EVSI was evaluated at every 0.01 point ($z$ values) between 0 and 1 (possible range for risks). Figure 2 illustrates computed EVSI for different sample sizes (coloured curves) at different threshold values, $z$, compared to the computed EVPI (black curve).
Figure 2: Computed EVSI (dotted/dashed curves) for future samples with four different seizes compared to the computed EVPI (solid curve).
EVPI quantifies the expected NB loss per decision (every time the model might be used). The overall NB loss due to uncertainty is therefore affected by the expected number of times the medical decision will be made at the target population (Willan and others, 2012). At the 0.21 threshold, a future validation study can have a maximum benefit equal to 0.003. One can put the VoI gain into context of the target population following the scheme proposed by Sadatsafavi and others (2023a). There are around 800,000 AMIs in the US every year (Virani and others, 2021). That is, at the 0.21 threshold value, a future validation study can have a maximum benefit equal to gaining 2050 net true positive (individuals who will die within 30 days and are correctly identified as high risk) per year, or avoiding 7709 extra false positive cases per year (individuals who will not die within 30 days).
As expected, the larger the future samples, the greater EVSI, and consequently, the closer curves to that of EVPI. The overall pattern of computed EVSIs remain similar to that of EVPI for different sample sizes across threshold values. The ‘diminishing return’ pattern is also obvious for both thresholds: the expected NB gain steeply rises at small samples and plateaus as $n$ grows. At 0.21 threshold, a future study of size $n=1000$ has a per-decision EVSI value of 0.00132, corresponding to population value of 1058 in true positive cases gained (or 3978 in false positive cases averted), while a future study of size $n=100000$ has a per-decision EVSI value of 0.00234, corresponding to population value of 1869 in true positive cases gained (or 7030 in false positive cases averted).
@@ 4 Discussion
In this study, we applied the value of information (VoI) framework to assess the decision-theoretic consequences of uncertainty due to the finite size of the development sample. We introduced the expected value of sample information (EVSI) for model development, defining it as the anticipated increase in net benefit (NB) gained from obtaining an additional development sample of a specified size from the target population. We proposed a bootstrap-based algorithm for calculating EVSI and demonstrated its feasibility through a case study. We further examined the relationship between EVSI and future development sample sizes, suggesting that scaling EVSI to the population level can quantify the total increase in clinical utility in terms of true or false positives.
Contemporary approaches for the assessment of uncertainty in the predictions of a model are rooted in classical inferential statistics. The incorporation of decision-theoretic approaches that quantify the expected gain by procuring further development data can provide a complementary perspective. Specifically, we suggest the computation of EVPI (at relevant thresholds) to quantify the expected utility loss due to uncertainty with the current development sample. If this loss was deemed high, the EVSI curve can be generated to evaluate the expected return-on-investment from a future development sample. The EVSI curve can be juxtaposed to multi-criteria sample size formulas for model development (Riley and others 2024). This might demonstrate, for example, if one of the specific criteria falls in the “diminishing return” part of the EVSI curve, and thus enables the investigator team to have an informed discussion about sample size requirements of a future validation study. As a case in point, we go back to our GUSTO-I case study. There, applying the equations by Riley and others (2024) at the threshold value of $z=0.02$ results in future sample size estimates of 159,170 for the observed/expected (O/E) criterion (targeting SE of 0.051), 2,477 for the C-slope criterion (targeting SE of 0.051), and 30,938 for the C-statistic criterion (targeting SE of 0.026). Notably, two out of three of these estimates, along with the final recommendation, fall within the “diminishing return” region ($n>10,000$), where further increasing the sample size would provide minimal benefit according to EVSI (Sadatsafavi and others, 2023b). Based on this, investigators might be able to justify settling on the sample sizes determined by the other two criteria.
@@
Recently, the impact of prediction uncertainty has been assessed in terms of the instability of predicted risks (Riley and Collins 2023). While prediction instability is indeed a relevant concerned, its impact on the overall utility of subsequent decisions is not clear. As well, to what extent a future development sample of a given size will reduce instability was not studied by the investigators. We note that the bootstrap-based approach for quantifying prediction instability, as proposed by Riley and Collins (2023) et al, is similar in operation to the bootstrap-based algorithms for VoI analysis. As such, one set of bootstrap-based simulations can be used for the evaluation of prediction instability, EVPI, and EVSI.
This research represents a continued effort to address prediction model uncertainty due to finite sample sizes (Sadatsafavi and others 2022; Sadatsafavi and others 2023a; Sadatsafavi and others 2023b). However, prediction models are also subject to other sources of uncertainty, such as measurement errors in predictors and outcomes or potential discrepancies between the development sample and the target population. Future research could explore the impact of other sources of uncertainty (e.g., whether the sample is truly representative of the population, whether future models will include additional covariates). Additionally, since EVSI itself is derived from a finite sample, it is subject to similar uncertainties as the prediction model. Assessing EVSI uncertainty, potentially by incorporating an additional level of bootstrapping into our EVSI proposed algorithm, could be another avenue for future work. Moreover, the present method assumes the availability of pilot data from the target population, and future studies should investigate how EVSI methods could be developed for broader contexts, including situations where no pilot data are available. This would enhance the integration of EVSI with sample size calculations (e.g., Riley and others 2024), potentially providing more flexible and generalizable approaches. Feature selection is another critical step in prediction model development, with various methods, such as LASSO, being used to refine models by regularizing and eliminating parameters based on a tuning parameter. Further research is needed to assess how model development processes, like feature selection, influence EVSI calculations. More importantly, EVSI could be incorporated into feature selection (generating EVSI values fro models of different complexity). This will help model developers in their quest to strive a balance between model accuracy and parsimony, guiding investigators to focus on collecting only the most relevant predictors in future model development. By doing so, EVSI can help reduce costs associated with prediction model development by tailoring both the number of observations and the number of predictors in future samples.

### v2

@@ 2.4 Expected value of sample information (EVSI)
Assume for the moment that we know the true case-mix ($\eta$) and the correct model ($\theta$). As such, we can generate data of the future study $D^{*}$. Given the assumption on the exchangability of $d$ and $D^{*}$, $D^{*}$ would be combined with the development sample $d$, which we denote by $D^{+}=(d,D^{*})$, and the investigator would revisit their decision for declaring the superior decision based on $W(D^{+})$. The ENB associated with a given $D^{+}$ is $ENB\left(W(D^{+});D^{+}\right)$.
However, as the future data have not been collected, we will need to take expectation with respect to the predictive distribution of $D^{*}$, i.e., expected of ENB of future sample information. Therefore, EVSI can be defined as the difference between this expectation and the ENB of the current information as follows
$EVSI=$ $E_{D^{+}}\left[ENB\left(W(D^{+});D^{+}\right)\right]-ENB\left(W(d);d\right)$ (9)
where the first term on right-hand side (RHS) is the expected of ENB from future sample for model development and the second term is the ENB under current information. EVSI is a scalar quantity in the same NB unit as the clinical utility of the model. The higher the EVSI, the higher the expected benefit from a sample.
EVSI is a non-decreasing function of the sample size $n^{*}$, and as $n^{*}$ increases, by the Law of Large Numbers, it converges to EVPI for model development, which can be expressed as
$EVPI=ENB_{max}(d)-ENB\left(W(d);d\right),$ (10)
where $ENB_{max}(d)$ is the expected of the estimated NB of the optimal strategy, which is based on using the true model itself for risk stratification
$NB_{max}\left(\eta,\theta\right)=\mathbb{E}\left\{I\left(\pi\left(\mathbf{X},\theta\right)\geq z\right)\left[\pi(\mathbf{X},\theta)-(1-\pi(\mathbf{X},\theta))\frac{z}{1-z}\right]\right\}.$ (11)
@@ 2.5 Bootstrap-based EVSI computation
The second term on the RHS of the EVSI formula in (9) is the ENB under current information and can be estimated following the same bootstrap algorithm given in Section 2.3. Therefore, to compute EVSI, it remains to estimate the first term. As the equation for EVSI implies, the nested computation of the two expectations generally gives rises to nested Monte Carlo algorithms (Ades and others, 2004; Kunst and others, 2020). Similarly, we will employ a nested Bayesian bootstrap algorithm. We note that, because the Bayesian bootstrap allows one to interpret $d^{t}$ as a random draw from the population that has generated $d$, we can use $d^{t}$ to also generate the sample for the planned future study. Specifically, we take a (second level of) bootstrap sample from $d^{(t)}$, called $d^{(*t)}=(\mathbf{x}^{(*t)},y^{(*t)})$, as our future sample of size $n^{*}$ given $\theta^{(t)}$. Note that, since we drew $d^{(*t)}$ under the assumption of $\pi^{(t)}$s being the true risks for the population individuals, we generate $Y^{(*t)}_{i}$s from Bernoulli distributions with their corresponding true risks being the success probabilities. Next, we update the original prediction risk model based on the pooled dataset of $d^{(+t)}=(d,d^{(*t)})$ as well as the predicted risk vector of $\pi\left(\mathbf{x}^{(+t)},\hat{\theta}^{(t)}\right)$ obtained from the updated model. The bootstrap-based EVSI computation algorithm presented in Algorithm 1.
@@
1:  for $t=1,\dots,T$ (number of simulations) do
2:     Randomly draw from the generating population $d^{(t)}$: Bayesian bootstrap sample from original (pilot) dataset. This is interpreted as a random draw from the distribution of the population.
3:     Fit true model: fit a new model on the bootstrapped sample $d^{(t)}$ and treat that as the true model
4:     Obtain true risks $\mathbf{\pi}^{(t)}$: estimate risks for patients in $d$ from the fitted model in Step 3
5:     Derive true NBs: apply true risks from step 3 in equation (5) to derive true NBs for patients in $d$
6:     Derive true NB of perfect information: apply true risks from step 3 in equation (11) to derive true NB of perfect information for patients in $d$
7:     Draw future sample $d^{(*t)}$: draw bootstrap sample of size $n^{*}$ from $\mathbf{x}^{(t)}$ of $d^{(t)}$
8:     Generate true outcomes: generate a new set of outcomes $Y^{(*t)}$ from Bernoulli distribution with the corresponding $\pi^{(t)}$
9:     Merge sample ($d^{(+t)}$): merge the future sample $d^{(*t)}$ with the development sample $d$
10:     Develop prediction model: develop a new prediction model using the merged data $d^{(+t)}$ and update the true risks accordingly
11:     Estimate NBs: estimate according to (5) on the merged sample
12:  end for
13:  Estimate ENBs: average estimated NBs based on $d$ according to (8)
14:  Estimate ENBs: average estimated NBs based on $D^{+}$s according to (8)
15:  Identify winner strategy: identify the strategy with maximum estimated ENB obtained from the previous step
16:  EVPI: Compute as the difference between ENB of perfect information and ENB of winner strategy
17:  EVSI: Compute according to (9)
@@ 3 Case study: GUSTO trial
As a case study, we apply these developments to GUSTO-I, a clinical trial of multiple thrombolytic strategies for acute myocardial infarction (AMI, or ‘heart attack’) (GUSTO Investigators, 1993). The GUSTO trial was a four-arm randomized clinical trial that compared the efficacy of different intravenous thrombolytic regimens for AMI. This dataset has been widely used for methodological research in predictive analytics (Ennis and others, 1998). We use these data to build a prediction model for predicting 30-day mortality for AMI. We used sex, age, diabetes (yes or no), milocc (which part of the heart muscle is affected by the blocked artery and coded as 3=anterior, 4=inferior, 5=other, and 6=no MI), previous MI, hypertension, smoking, killip class, and treatment as covariates in the model. Similar to our previous work, we set a risk threshold of 1% as the default threshold, but will evaluate the model at the entire possible range of thresholds.
Given the relatively large sample size (n=40,830), the uncertainty in predictions when the model is fitted using the full sample is negligible. Indeed, a previous study reported an EVPI of 0 with the full dataset (thus, EVSI will be zero for all sample sizes). Therefore, we took, without replacement, a random sub-sample of 1,000 patients from GUSTO data and treat that as our development sample (source of current information). The logistic regression model fitted on this sample is given Appendix 6.
We start by drawing the decision curve at the entire possible range of thresholds, as shown in Figure 1 (panel A). The figure shows the NB of the model as well as the alternative default strategies (treat none and treat all). The model had higher expected NB than the alternative strategies at the all threshold values of interest. Additionally, panel (B) of Figure 1 shows the difference between the estimated NB of the model and that of the best alternative default strategy, i.e., $ENB(1,d)-max\left(0,ENB(2,d)\right)$, along with its bootstrap $95\%$ CI. Again, the model had a higher estimated NB than the default alternatives at the entire range of threshold values of interest.
@@
Next, we draw EVSI. We chose a range of future sample size, spanning from 100 to 32,000. Following algorithm 1, for each sample size, we drew $10^{5}$ bootstrap samples ($T=10^{5}$) from the development sample, and EVSI was evaluated at every 0.01 point ($z$ values) between 0 and 1. Panel A of Figure 2 illustrates computed EVSI (right Y-axis) for various sample sizes (dashed and dotted lines) at different threshold values, $z$, alongside the computed EVPI (black curve). Both EVPI and EVSI quantify the expected NB loss per decision each time the model is applied. The overall NB loss attributable to uncertainty is influenced by the expected frequency of the medical decision being made (Willan and others, 2012). To scale the VoI metrics (EVPI and EVSI) to the population level, given that approximately 800,000 acute myocardial infarctions (AMIs) occur annually in the United States (Virani and others, 2021), we provide scaled VoI values in terms of true positive units on the right Y-axis of Panel A in Figure 2. At a threshold of 0.01, where the EVPI is 0.0001, a future development study could yield a maximum benefit equivalent to identifying 89 additional true positive cases (individuals who will die within 30 days and are correctly classified as high risk) per year, or avoiding 8,761 false positive cases (individuals who will not die within 30 days but are incorrectly classified as high risk) annually. Panel B Figure 2 shows the EVSI curve at threshold value of $z=0.01$. As anticipated, EVSI increases with larger future sample sizes, eventually asymptoting to the EVPI. The expected NB gain (left Y-axis) exhibits a steep increase at smaller sample sizes, followed by a plateau as $n^{*}$ grows, reflecting a “diminishing return” pattern. Similar to the EVPI plot in Panel A, we scale the EVSI to the population level, with the corresponding values displayed on the right Y-axis of Panel B in Figure 2. At 0.01 threshold, a future study of size $n^{*}=1,000$ has a per-decision EVSI value of 0.000049, corresponding to population value of 39 in true positive cases gained per year (or 3,881 in false positive cases averted annually), while a future study of size $n^{*}=16,000$ has a per-decision EVSI value of 0.00010, corresponding to population value of 82 in true positive cases gained per year (or 8,141 in false positive cases averted annually).
Figure 2: Panel (A): Computed EVSI (dotted/dashed curves) for future samples with different sizes compared to the computed EVPI (solid curve) at different threshold values in terms of N difference (left Y-axis) and EVPI scaling at population (right Y-axis). Panel (B): Computed EVSI (left Y-axis) and its scaling at population (right Y-axis) with different future sample sizes (red curve) at threshold value of $z=0.01$ compared to its corresponding computed EVPI (black curve). Dotted/dashed vertical curves are the estimated sample sizes based on the three criteria of Riley and others (2024): observed/expected ratio (Criterion 1 - dashed), calibration slope (Criterion 2 - dotted), and C-statistic (Criterion 3 - dashed-dotted).
For comparison, we provide sample size calculations results according to Riley and others (2024)’s criteria. Panel B of Figure 2 displays vertical dashed/dotted lines indicating the estimated sample sizes required to develop a new multivariable prediction model, calculated using the three criteria proposed by Riley and others (2024). These estimates were obtained using the $pmsampsize$ function from the R package $pmsampsize$ (version 1.1.3; Ensor, 2023). We specified the following inputs based on the development sample: a binary outcome, a C-statistic of 0.746, 13 candidate predictors, and an outcome prevalence of 0.074. The resulting sample size estimates were: 2,013 for the observed/expected (O/E) ratio (Criterion 1), 610 for the calibration slope (Criterion 2), and 106 for the C-statistic (Criterion 3). Following Riley and others (2024), the largest of these values (2,013) was selected as the recommended sample size for model development. The expected NB gain changes by $\sim 50\%$ when moving from Criterion 1 to 2 and $\sim 75\%$ when moving from 2 to 3.
@@ 4 Discussion
In this study, we applied the VoI framework to assess the decision-theoretic consequences of uncertainty due to the finite size of the development sample. We introduced the expected value of sample information (EVSI) for model development, defining it as the anticipated increase in net benefit (NB) gained from obtaining an additional development sample of a specified size from the target population. We proposed a bootstrap-based algorithm for calculating EVSI and demonstrated its feasibility through a case study. We further examined the relationship between EVSI and future development sample sizes, suggesting that scaling EVSI to the population level can quantify the total increase in clinical utility in terms of true or false positives.
A practical application for this study is at the end of a model development study, by calculating EVPI and EVSI, one can decide whether they should focus on obtaining more development sample, or move on to the next stage (e.g., independent validation or an impact study). Specifically, we suggest the computation of EVPI (at relevant thresholds) to quantify the expected utility loss due to uncertainty with the current development sample. If this loss was deemed high, the EVSI curve can be generated to evaluate the expected return-on-investment from a future development sample. The EVSI curve can be juxtaposed to multi-criteria sample size formulas for model development (Riley and others 2024). By comparing these criteria against the EVSI curve, researchers can assess whether proposed sample sizes fall within the steeply increasing phase (where marginal gains are substantial) or the plateau region (reflecting diminishing returns). In our GUSTO-I case study, we applied Riley and others (2024)’s criteria to estimate the required sample size for a development study. Notably, all three estimates, along with the final recommended sample size of 2,013, fell within the steeply ascending portion of the EVSI curve (where $n<4,000$ for $z=0.01$). This suggests that increasing the sample size beyond these estimates would still yield substantial improvements in net benefit, as quantified by EVSI. Consequently, investigators may need to reevaluate whether conventional sample size criteria adequately account for the marginal value of additional data, particularly when the EVSI curve indicates substantial unrealized gains.
@@
Recently, the impact of prediction uncertainty has been assessed in terms of the instability of predicted risks (Riley and Collins 2023). While prediction instability is indeed a relevant concern, its impact on the overall utility of subsequent decisions is not clear. As well, to what extent a future development sample of a given size will reduce instability was not studied by the investigators. We note that the bootstrap-based approach for quantifying prediction instability, as proposed by Riley and Collins (2023) et al, is similar in operation to the bootstrap-based algorithms for VoI analysis. As such, one set of bootstrap-based simulations can be used for the evaluation of prediction instability, EVPI, and EVSI.
The current method assumes data from the target population are already available, and thus is best suited for the situation where either pilot data are at hand, or after the model is developed. However, in many cases, developers lack such data for a new population and must rely on evidence from similar models in other populations. Specifying uncertainty around model performance based on previous studies and expert opinion, and developing VoI algorithms that do not rely on an existing sample, is the natural next step. Riley and others (2025) recently proposed an approach for sample size calculation for development studies based on decomposing the variance of an individual’s risk estimate into components based on Fisher’s unit information matrix, predictor values, and total sample size. This method shows promise for informing VoI analyses. Further, prediction models are also subject to other sources of uncertainty, such as measurement errors in predictors and outcomes or potential discrepancies between the development sample and the target population. Future research could explore the impact of other sources of uncertainty (e.g., whether the sample is truly representative of the population, whether future models will include additional covariates). Additionally, EVSI is computed using Monte Carlo algorithms, and thus itself is associated with uncertainty due to Monte Carlo error. Objectively deciding on the number of Monte Carlo simulations can be pursued in future research. Feature selection is another critical step in prediction model development, with various methods, such as uniLASSO (Chatterjee and others, 2025), being used to refine models by regularizing and eliminating parameters based on a tuning parameter. We note that the Bayesian interpretation of shrinkage methods (Mitchell and Beauchamp, 1988) justifies their use on our bootstrapped-based algorithm. In particular, shrinkage can be applied when generating the correct model (Step 3 of Algorithm 1), incorporating our prior belief that coefficients cannot take extremely high values. Further research is needed to assess how model development processes, like feature selection, influence EVSI calculations.
@@
5 Software
Software in the form of R code, together with a sample input data set and complete documentation is available on our GitHub page at https://github.com/ab-sa/evsi_modDev.
6 Apendix: GUSTO prediction model
In the case study, we used a sub-sample of 1000 patients from GUSTO data and treat that as our development sample. Using the same development sample, we fitted a logistic regression model and treat that as the current prediction model on the current sample. The fitted model along with the estimated coefficient was as follows:
$logit(mortality)$ $=-6.56~{}Sex_{male}+-6.26~{}Sex_{female}+0.06~{}Age-0.02~{}Diabetes+0.53~{}milocc_{Other}$
$+0.11~{}milocc_{Anterior}+0.72~{}Prev_{MI}+0.05~{}Hypertension-0.41~{}Smoking_{quit}$
$-0.35~{}Smoking_{current}+1.02~{}Killip-0.27~{}Treatment_{PA}$